Of the first 15 sandwiches sold in one day, 9 were club sandwiches. At the same rate, how many sandwiches would have to be sold for 90 to be club sandwiches.
Answer by TimothyLamb(4379)

total/club = 15/9
total/90 = 15/9
total = 90*15/9
90*15/9
---
copy and paste the above expression in to this calculator:
http://sooeet.com/math/scientific-calculator.php
---
total = 150
---
Free algebra tutoring live chat:
http://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
http://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
http://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
http://sooeet.com/math/system-of-linear-equations-solver.php
---